5x^2+5y^2-20=0 in standard form
Answer by lwsshak3(11628)

5x^2+5y^2-20=0 in standard form
**
5x^2+5y^2-20=0
divide by 5
5x^2+5y^2-20=0
x^2+y^2-4=0
x^2+y^2=4
This is an equation of a circle with center at (0,0) and radius=2, of the standard form:
(x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center and r=radius